OpenStudy (anonymous):

What is the slope of (2,7) and (-6,21)?

OpenStudy (anonymous):

DO YOU KNOW THE FORMULA FOR THE SLOPE?

OpenStudy (anonymous):

is it like y2-y1 over x2 exc. that jon?

OpenStudy (anonymous):

YES M=Y2-Y1 / X2-X1

OpenStudy (anonymous):

slope = -7/4

OpenStudy (anonymous):

if we label the pairs that u are giving me we will have 2=x1 7=y1 -6=x2 21=y2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(2,7) and (-6,21) m= change in y/change in x m= (21-7)/(-6-2) m=14/8 m=7/4